Bartosiewicz Clinic uses client-visits as its measure of activity. During January, the clinic budgeted for 4,500 client-visits, but its actual level of activity was 4,480 client-visits. The clinic has provided the following data concerning the formulas used in its budgeting and its actual results for January: Data used in budgeting: Fixed element per month Variable element per client-visit Revenue - $ 36.50 Personnel expenses $ 28,500 $ 12.50 Medical supplies 1,400 6.50 Occupancy expenses 10,400 1.00 Administrative expenses 7,500 0.10 Total expenses $ 47,800 $ 20.10 Actual results for January: Revenue $ 129,000 Personnel expenses $ 70,010 Medical supplies $ 21,010 Occupancy expenses $ 13,180 Administrative expenses $ 7,430 The overall revenue and spending variance (i.e., the variance for net operating income in the revenue and spending variance column on the flexible budget performance report) for January would be closest to: Garrison 16e Rechecks 2018-08-12 Multiple Choice $8,630 F $8,630 U $8,302 F $8,302 U
Answer #1

Net income in Flexible budget = (36.50-20.10)*4480-47800 = 25672

Actual net operating income = 17370

Overall revenue and spending variance = Flexible budget-actual net operating income

= 25672-17370

Overall revenue and spending variance = 8302 U

So answer is d) $8302 U